Circular No. 6036

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) T. Encrenaz, R. Schulz, J. A. Stuewe and G. Wiedemann communicate:

“We report the detection of near-infrared emissions in the H_{2} S(1) line (2.12 μm) and in the 3.533-μm H_{3}^{+} line in the northern hemisphere of Jupiter at latitude +44^{o} and at the longitudes of the brightest impacts regions. These emissions have been detected with the IRSPEC imaging spectrometer at the 3.5-m NTT at the European Southern Observatory. On July 23.017–23.073 UT the northern counterpart of the R-S-G-Q region was detected both in the H_{2} and the H_{3}^{+} lines. The H_{2} line, of the same intensity as in the R-S-G-Q region, was superimposed over a continuum three times weaker than in the corresponding impact region. The H_{3}^{+} line was also easily detected, over a weak continuum. Later (July 23.17 UT), the northern counterparts of sites A and V-T-E-F were also detected. These emissions disappeared at lower northern latitude. The presence of emissions at latitude +44^{o} was again confirmed on July 23.83 UT, with the H_{3}^{+} detection of the northern counterparts of impacts R-S, Q, H and V-T-E-F. At latitude +44^{o} the H_{3}^{+} emission is spread over the entire longitude range, with strong maxima in the northern counterparts of the impact sites. There is little continuum as compared to the 2.12-μm emission. As suggested by several other observers, these emissions seem to be the signature of particle transfer along the lines of the jovian magnetic field.”

COMET NAKAMURA-NISHIMURA-MACHHOLZ (1994m) Improved parabolic orbital elements from MPC 23766:

T = 1994 July 12.902 TT  ω = 123.^{o}005
Ω = 158.960 }2000.0
q = 1.14088 AU  i = 94.388
1994 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
July 27 2^{h}41.^{m}11 +68^{o}55.^{′}1 0.998 1.163 70.^{o}6 55.^{o}4 8.7
Aug. 1 2 12.44 +67 55.0 0.894 1.181  76.2 56.5 8.5
6 1 38.23 +66 10.6 0.787 1.204  82.9 56.7 8.3
11 0 58.91 +63 09.2 0.682 1.232  91.1 55.3 8.1
16 0 16.48 +57 59.2 0.582 1.264 101.4 51.7 7.8
21 23 34.31 +49 28.5 0.494 1.300 114.9 44.9 7.6
26 22 55.81 +36 28.6 0.431 1.338 132.2 34.0 7.4
31 22 23.13 +19 24.3 0.406 1.380 151.6 20.3 7.4
Sept. 5 21 56.90 + 1 40.6 0.429 1.425 163.2 11.8 7.7

1994 July 24 Brian G. Marsden